OpenStudy (anonymous):

Rewrite as a sum or difference ln(ex^3)

OpenStudy (anonymous):

\[\ln (ab) = \ln a + \ln b\] also note that \[\ln e = 1\]

OpenStudy (anonymous):

One more thing, @abb50, \[\ln(u^x)=x\ln(u).\]

OpenStudy (anonymous):

ah yes of course. thank you for reminding